# Bridging the Linguistic Divide: A Survey on Leveraging Large Language Models for Machine Translation

### v2

@@ 1 Introduction
Machine Translation (MT) has significantly evolved from early rule-based systems to advanced neural models. Initially, MT relied heavily on manually crafted linguistic rules and dictionaries, requiring extensive human intervention for each language pair and struggling with linguistic nuances, idiomatic expressions, and contextual understanding [1]. The transition to Statistical Machine Translation (SMT) in the late 1980s and early 1990s marked a major advancement, as SMT utilized probabilistic models trained on large bilingual corpora, automating translation processes [2, 3]. However, SMT faced challenges such as limited fluency, difficulties with less common language pairs, and issues handling long-range dependencies. The introduction of neural networks marked a turning point in MT research. Recurrent Neural Networks (RNNs) and later, sequence-to-sequence models with Long Short-Term Memory (LSTM) [4] units improved translation fluency by learning contextual dependencies [5, 6]. However, these architectures faced limitations in handling long sequences due to vanishing gradients and computational constraints. The emergence of attention mechanisms [7, 8] addressed these issues by allowing models to focus selectively on relevant parts of the input sequence, enhancing both accuracy and coherence. This led to the development of the Transformer model [9], which became the foundation for modern neural MT systems.
@@
Large Language Models (LLMs) offer new opportunities to address these challenges. Their ability to learn from vast amounts of multilingual and monolingual text allows them to generate translations with improved fluency and contextual awareness. Few-shot and zero-shot learning capabilities enable LLMs to handle low-resource languages with minimal supervision. Moreover, LLMs facilitate domain adaptation and personalized translation by leveraging in-context learning. However, they also introduce concerns such as increased computational costs, potential biases, and the necessity for efficient deployment strategies.
As MT continues to advance, the integration of LLMs is expected to redefine the field, offering new possibilities while necessitating careful consideration of ethical and computational trade-offs. Future research must focus on improving efficiency, addressing bias, and ensuring reliable translations across diverse languages and domains.
@@ 2 Structure of the Survey
This survey provides a comprehensive analysis of the role of Large Language Models (LLMs) in Machine Translation (MT), particularly focusing on low-resource languages. The paper is structured as follows:
@@
This study offers a comprehensive analysis of how large language models (LLMs) are transforming machine translation (MT), particularly in low-resource settings where parallel data is limited. It highlights the emerging challenges and opportunities, providing a foundation for future research in this rapidly evolving domain. To ensure methodological consistency, the study refrains from direct cross-paper comparisons, as existing works vary significantly in terms of datasets, language pairs, and evaluation metrics.
For clarity and focus, the analysis is restricted to text-to-text, sentence-level translation, excluding document-level, dialogue-based, and multimodal approaches (e.g., image or video-assisted translation, sign language translation). While certain studies employ hybrid techniques—such as reinforcement learning (RL) combined with quantized low-rank adaptation (QLoRA)—they are categorized under the primary methodology most representative of their core contribution.
@@ 3 Synthetic Data Generation with LLMs
Domain adaptation in Machine Translation (MT) has been explored by leveraging large pre-trained language models (LMs) for synthetic data generation [26]. This approach addresses two scenarios: one with limited bilingual in-domain data and another with no bilingual in-domain data. In the former case, target-side synthetic text is generated using GPT-J (English) and mGPT (Arabic), followed by back-translation to create parallel data. In the latter case, source text is forward-translated using an MT model before applying the same augmentation technique. A mixed fine-tuning strategy with in-domain data oversampling is employed to ensure balanced learning, resulting in significant BLEU [27, 28] score improvements of approximately 5-6 for Arabic-English and 2-3 for English-Arabic. The improvement in quality has been also confirmed by manual evaluations.
@@
The use of word and phrase alignment for synthetic data generation has also been studied [30]. However, findings indicate that synthetic data produced by LLMs does not offer significant improvements over rule-based and neural-based approaches [31].
@@ 4 Translation Methodologies
Machine Translation using LLMs can be broadly categorized into two primary methodological approaches: (1) prompting-based techniques and (2) fine-tuning-based techniques. The former leverages pre-trained models with minimal adaptation, whereas the latter involves modifying model parameters to enhance performance on translation tasks. The methods can be combined - prompting based techniques can be used after fine-tuning. In this section, we discuss these methodologies in detail.
@@ 4.1.1 Zero-Shot Prompting
Different approaches in zero-shot prompting vary by prompt format and the use of a pivot language for low-resource languages. It has been observed that zero-shot prompting with ChatGPT lags behind MT systems by Google MT [32], Tencent [33], and DeepL by around 5.0 BLEU points [34]. Pivot prompting has been explored to translate between distant languages, where the LLM first translates the sentence to English and then into the target language. This strategy, which uses a resource-rich language (English) as a pivot, improves translation quality between DeZh and RoZh [34]. Human evaluation on the generated translations reveals that GPT-4 makes fewer mistakes compared to ChatGPT and Google MT. Interestingly, although GPT-4 scores lower in BLEU compared to Google MT, human annotators rank its translations more favorably in blind evaluations [34].
@@ 4.1.2 In-context Learning
In-Context Learning (ICL) enhances machine translation by providing a model with a small set of input-output examples before generating a translation, helping it infer correct patterns. This method significantly improves translation accuracy, especially for domain-specific or idiomatic expressions, by guiding the model toward appropriate mappings. The effectiveness of the approach depends on multiple parameters, including the number of in-context examples, selection of the examples, and their ordering.
@@
Number of In-Context Examples: One critical decision when designing prompts for large language models (LLMs) is determining the optimal number of in-context examples. Increasing the number of examples can enhance performance, but it also comes with higher computational costs and diminishing returns. Using too few examples may bias the model towards the style of the provided instances, while incorporating too many may not lead to proportional improvements in output quality.
@@
Further advancing example selection strategies, research has highlighted the importance of coherence in in-context learning [48]. Rather than relying solely on similarity-based retrieval, coherence is defined as the semantic and stylistic consistency between prompt examples and the test sentence. Studies show that ensuring coherence—aligning prompts in meaning, style, and contextual flow—yields better translation performance than selecting examples purely based on word or embedding similarity. This approach shifts the focus from isolated, highly similar examples to logically consistent prompt setups, redefining how in-context examples should be chosen for machine translation.
@@
The authors in [42] used translation memories (TMs) to enhance large language model (LLM) translation performance. The approach retrieves the top-k most similar sentence pairs from a TM database and incorporates them into prompts using structured templates. Two main template styles—instruction-style and code-style—are used to provide translation hints. Experiments demonstrate that TM-based prompting significantly improves LLM translation quality, often surpassing strong NMT baselines.
@@ 4.2.1 Supervised Fine-Tuning (SFT)
Supervised fine-tuning (SFT) involves training an LLM on a large parallel corpus by minimizing the cross-entropy loss between predicted translations and ground truth references. This method significantly enhances the model’s translation accuracy, especially for specialized domains. It has been shown that SFT enables LLMs to achieve strong multilingual translation performance with as few as 32 examples, where training on one language direction generalizes to others [55]. However, using English as the target language can lead to misinterpretations, and the presence of target-side noise tends to affect well-represented languages more adversely. Ensuring successful SFT alignment requires the model to maintain a surface-level focus, preventing it from acquiring unwanted biases during the fine-tuning process.
@@ 4.2.2 Low-Rank Adaptation (LoRA)
LoRA [56] is a parameter-efficient fine-tuning technique that reduces memory overhead by injecting small trainable matrices into frozen pre-trained models. This approach enables effective adaptation to new translation domains without modifying the entire model. A comprehensive empirical study on parameter-efficient fine-tuning (PEFT) methods for machine translation, conducted across various parameter budgets and language pairs, observed that using adapters achieves performance on par with full fine-tuning while training only 10% of the parameters [57]. However, the effectiveness of PEFT methods varies significantly based on language pair distance and dataset size, with noticeable struggles for distant language pairs. We refer the readers to [58] and [59] for a detailed analysis for PEFT methods.
@@
An evaluation of 15 publicly available language models on machine translation tasks showed that QLoRA fine-tuning outperforms both few-shot learning and models trained from scratch, achieving a significant BLEU score improvement while training on only 0.77% of the model parameters [63]. Adapter-based fine-tuning with LoRA was further shown to match the performance of traditional fine-tuning, while reducing the number of training parameters by a factor of 50 [39].
4.2.3 ALMA
A two-stage fine-tuning approach has been introduced to enhance machine translation capabilities in large language models (LLMs) [64]. Since many LLMs are predominantly trained on English data, the first stage, known as ALMA, fine-tunes the model on non-English monolingual datasets. This step strengthens the model’s proficiency in various languages, enabling a more robust understanding of linguistic nuances across different tongues.
@@ 4.2.4 Reinforcement Learning based Finetuning
The standard SFT methodology, which focuses on reducing the gap between predicted outputs and gold-standard references, inherently limits model performance to the quality of the training data. This presents a notable challenge, as even human-authored data—typically regarded as high-quality—can contain imperfections. To address this limitation, the authors in [65] introduced Contrastive Preference Optimization (CPO) for machine translation. CPO differentiates itself by training models to distinguish between “adequate” and “near-perfect” translations. This is achieved through a contrastive learning strategy that presents the model with hard negative examples—translations that are high-quality but not flawless. By learning to reject these near-perfect yet imperfect translations, the model is encouraged to produce superior outputs. When applied to ALMA models with only 22K parallel sentences and fine-tuning only 0.1% of the parameters, CPO yielded significant improvements. The authors in [66] uses Reinforcement Learning with Human Feedback (RLHF) to improve machine translation by aligning model outputs with human preferences. Instead of relying on expensive human-annotated preference data, the authors use high-quality human translations from books as a proxy for preferred outputs. The process involves fine-tuning a base translation model on parallel corpora, training a reward model to distinguish between machine and human translations, and then refining the translation model through RL using Proximal Policy Optimization (PPO) [67]. This approach enhances translation quality and shows cross-lingual transferability, meaning the improvements generalize to languages not directly optimized with RLHF. The results demonstrate that RLHF can effectively capture human translation preferences in a cost-efficient manner.
5 LLMs vs Encoder-Decoder across Languages

### v3

@@ 1 Introduction
Machine Translation (MT) has significantly evolved from early rule-based systems to advanced neural models. Initially, MT relied heavily on manually crafted linguistic rules and dictionaries, requiring extensive human intervention for each language pair and struggling with linguistic nuances, idiomatic expressions, and contextual understanding [1]. The transition to Statistical MT (SMT) in the late 1980s and early 1990s marked a major advancement, as SMT utilized probabilistic models trained on large bilingual corpora, automating translation processes [2, 3]. However, SMT faced challenges such as limited fluency, difficulties with less common language pairs, and issues handling long-range dependencies. The introduction of neural networks marked a turning point in MT research. Recurrent Neural Networks (RNNs) and later, sequence-to-sequence models with Long Short-Term Memory (LSTM) [4] units improved translation fluency by learning contextual dependencies [5, 6]. However, these architectures faced limitations in handling long sequences due to vanishing gradients and computational constraints. The emergence of attention mechanisms [7, 8] addressed these issues by allowing models to focus selectively on relevant parts of the input sequence, enhancing both accuracy and coherence. This led to the development of the Transformer model [9], which became the foundation for modern neural MT systems. MT is now widely integrated into global communication—from business collaborations to cross-cultural teamwork enabling multilingual interactions in real time or asynchronously. A study [10] show that MT-mediated exchanges significantly enhance communication clarity, discussion depth, and decision-making quality in multilingual team settings.
@@
LLMs offer new opportunities to address these challenges. Their ability to learn from vast amounts of multilingual and monolingual text allows them to generate translations with improved fluency and contextual awareness. Few-shot and zero-shot learning capabilities enable LLMs to handle low-resource languages with minimal supervision. Moreover, LLMs facilitate domain adaptation and personalized translation by leveraging in-context learning. However, they also introduce concerns such as increased computational costs, potential biases, and the necessity for efficient deployment strategies.
As MT continues to evolve, the integration of LLMs is poised to fundamentally transform the field—introducing new capabilities while also raising important ethical and computational considerations. Although prior surveys have reviewed various aspects of MT [27, 28, 29], to the best of our knowledge, none have systematically captured the substantial advancements enabled by LLMs in recent years. Given the diverse array of techniques available—ranging from prompting strategies to parameter-efficient fine-tuning—navigating this emerging landscape can be challenging for newcomers and practitioners alike. To address this gap, this paper presents a comprehensive analysis of the principal methodologies for leveraging LLMs in MT, offering detailed comparisons and insights to guide future research and application.
@@ 2 Structure of the Survey
This survey provides a comprehensive analysis of the role of LLM in MT, particularly focusing on low-resource languages. The paper is structured as follows:
@@
This study offers a comprehensive analysis of how LLMs are being used for MT. It highlights the emerging challenges and opportunities, providing a foundation for future research in this rapidly evolving domain. To ensure methodological consistency, the study refrains from direct cross-paper comparisons, as existing works vary significantly in terms of datasets, language pairs, and evaluation metrics.
For clarity and focus, the analysis is restricted to text-to-text, sentence-level translation, excluding document-level[30, 31], dialogue-based, and multimodal approaches [32] (e.g., image or video-assisted translation, sign language translation [33]). While certain studies employ hybrid techniques—such as reinforcement learning (RL) combined with quantized low-rank adaptation (QLoRA)—they are categorized under the primary methodology most representative of their core contribution.
@@ 3 Synthetic Data Generation with LLMs
Domain adaptation in MT has been explored by leveraging pre-trained LLMs for synthetic data generation [34]. This approach addresses two scenarios: one with limited bilingual in-domain data and another with no bilingual in-domain data. In the former case, target-side synthetic text is generated using GPT-J [35] (English) and mGPT [36] (Arabic), followed by back-translation [37] to create parallel data. In the latter case, source text is forward-translated using an MT model before applying the same augmentation technique. A mixed fine-tuning strategy with in-domain data oversampling is employed to ensure balanced learning, resulting in significant BLEU [38, 39] score improvements of approximately 5-6 for Arabic-English and 2-3 for English-Arabic. The improvement in quality has been also confirmed by manual evaluations.
@@
The use of word and phrase alignment for synthetic data generation has also been studied [41]. However, findings indicate that synthetic data produced by LLMs does not offer significant improvements over rule-based and neural-based approaches [42], although the study focused on only one language. On a broader study, covering more languages, the authors in [43] generate a document-level synthetic corpus by forward-translating English Europarl data into seven diverse low-resource languages, later expanding it to 147 additional language pairs via pivoting. They evaluate the quality of this data using automatic metrics and human annotation, confirming its general reliability. Crucially, they show that MT models trained solely on this synthetic data can rival or outperform baselines trained on real data, and that fine-tuning strong pretrained systems like NLLB [44] and LLaMA-3B [45] on this data yields significant performance improvements. Furthermore, their experiments highlight that even noisy synthetic data can be highly effective when clean alternatives are scarce.
@@ 4 Translation Methodologies
MT using LLMs can be broadly categorized into two primary methodological approaches: (1) prompting-based techniques and (2) fine-tuning-based techniques. The former leverages pre-trained models with minimal adaptation, whereas the latter involves modifying model parameters to enhance performance on translation tasks. The methods can be combined - prompting based techniques can be used after fine-tuning. In this section, we discuss these methodologies in detail.
@@ 4.1.1 Zero-Shot Prompting
Different approaches in zero-shot prompting vary by prompt format and the use of a pivot language for low-resource languages. It has been observed that zero-shot prompting with ChatGPT lags behind MT systems by Google MT [46], Tencent [47], and DeepL by around 5.0 BLEU points [48]. Pivot prompting has been explored to translate between distant languages, where the LLM first translates the sentence to English and then into the target language. This strategy, which uses a resource-rich language (English) as a pivot, improves translation quality between DeZh and RoZh [48]. Human evaluation on the generated translations reveals that GPT-4 [49] makes fewer mistakes compared to ChatGPT and Google MT. Interestingly, although GPT-4 scores lower in BLEU compared to Google MT, human annotators rank its translations more favorably in blind evaluations [48].
@@ 4.1.2 In-context Learning
In-Context Learning (ICL) enhances MT by providing a model with a small set of input-output examples before generating a translation, helping it infer correct patterns. This method significantly improves translation accuracy, especially for domain-specific or idiomatic expressions, by guiding the model toward appropriate mappings. The effectiveness of the approach depends on multiple parameters, including the number of in-context examples, selection of the examples, and their ordering.
@@
Number of In-Context Examples: One critical decision when designing prompts for LLMs is determining the optimal number of in-context examples. Increasing the number of examples can enhance performance, but it also comes with higher computational costs and diminishing returns. Using too few examples may bias the model towards the style of the provided instances, while incorporating too many may not lead to proportional improvements in output quality.
@@
Further advancing example selection strategies, research has highlighted the importance of coherence in in-context learning [64]. Rather than relying solely on similarity-based retrieval, coherence is defined as the semantic and stylistic consistency between prompt examples and the test sentence. Studies show that ensuring coherence—aligning prompts in meaning, style, and contextual flow—yields better translation performance than selecting examples purely based on word or embedding similarity. This approach shifts the focus from isolated, highly similar examples to logically consistent prompt setups, redefining how in-context examples should be chosen for MT.
@@
The authors in [57] used translation memories (TMs) to enhance LLMs translation performance. The approach retrieves the top-k most similar sentence pairs from a TM database and incorporates them into prompts using structured templates. Two main template styles—instruction-style and code-style—are used to provide translation hints. Experiments demonstrate that TM-based prompting significantly improves LLM translation quality, often surpassing strong NMT baselines.
@@
4.2 Advanced Prompting
From a Chain-of-Thought (CoT) perspective, one of the early studies investigated whether step-by-step reasoning prompts, successful in many NLP tasks, can enhance ChatGPTś performance in MT [71]. They apply both zero-shot and 1-shot CoT prompting, where ChatGPT is instructed to translate source sentences word-by-word before generating the full translation. However, unlike in reasoning tasks, CoT induces a literal, phrase-by-phrase translation behavior that hampers fluency and semantic coherence, leading to a significant drop in COMET scores (e.g., -8.8 for ENZH). This degradation highlights a misalignment between CoT prompting and the holistic nature of translation, suggesting that while CoT can benefit structured reasoning, it is ill-suited for MT directly.
[…]
The MAPS (Multi-Aspect Prompting and Selection) [74] framework introduces a prompting strategy that mimics human translation behavior by guiding LLMs through a structured, three-step process. First, the LLM is prompted to extract translation-relevant knowledge from the source sentence in three forms—keywords, the underlying topic, and relevant example translations—using manually designed few-shot prompts. Next, each type of extracted knowledge is separately reintegrated into new prompts to guide the LLM in generating multiple translation candidates. Finally, a selection mechanism, either via a reference-free quality estimation model (COMET-QE) or prompting the LLM itself (LLM-SCQ), is used to choose the best candidate. This multi-faceted prompting strategy improves translation quality across various language pairs and also helps to reduce common errors such as hallucination, ambiguity, and mistranslation without relying on external data or curated in-context examples.
Another method which use Recognizing that translations generated by LLMs often contain subtle errors despite strong benchmark performance, the authors in [75] propose TEaR - a plug-and-play self-refinement framework for MT that improves translation quality through internal feedback loops. TEaR operates in three stages: Translate, where an LLM produces an initial translation; Estimate, where the same LLM simulates human-like error assessment using MQM guidelines; and Refine, where the LLM revises the translation based on its self-generated evaluation. This self-contained design allows TEaR to bypass reliance on external models or human feedback. The Estimate module identifies major and minor translation errors—such as mistranslation, awkward style, and grammar issues—and only proceeds to refinement if necessary, making the approach more selective and stable than baselines like CoT. Experiments on WMT22 and WMT23 datasets across eight translation directions show that TEaR consistently improves translation quality and reducing MQM error counts more effectively than other self-refinement or external feedback systems. Moreover, iterative application of TEaR further enhances translation quality in successive rounds, demonstrating its capacity for continuous improvement.
While prior works attempt to mimic human cognitive behaviors in prompting—such as self-reflection, iterative refinement, or chain-of-thought reasoning—to improve model performance, these methods often fall short when it comes to generating truly divergent or corrective reasoning paths. In particular, self-reflection-based approaches can suffer from what the authors term the Degeneration-of-Thought (DoT) problem: once an LLM gains confidence in its initial (possibly incorrect) answer, it becomes resistant to revising or challenging that stance in subsequent iterations. To address this, the paper proposes the Multi-Agent Debate (MAD) framework [76], which draws on the human problem-solving strategy of adversarial discussion. MAD simulates a structured debate between multiple LLM agents—typically an affirmative and a negative debater—who challenge each other in a tit for tat style, with a judge model overseeing the debate and determining the final answer.
This debate setup introduces external feedback and perspective shifts that self-reflection lacks, promoting more diverse and accurate reasoning. The MAD framework is evaluated on two challenging tasks—Commonsense MT and Counter-Intuitive Arithmetic Reasoning—where it significantly outperforms baseline methods including self-reflection and even matches or surpasses stronger models like GPT-4 when used with GPT-3.5 as the base. Key to MAD’s effectiveness are its adaptive break strategy (which ends debates early when a satisfactory solution is found), balanced argumentative intensity, and consistent use of the same LLM for all roles to mitigate judge bias. This methodology not only encourages divergent thinking in LLMs but also offers a scalable, model-agnostic approach to overcoming reasoning limitations in existing prompting strategies. Compositional Translation (CompTra) [77] is a prompting strategy designed to enhance the MT capabilities of LLMs, particularly for low-resource languages. Instead of translating entire sentences directly, CompTra treats translation as a step-by-step reasoning task. First, the input sentence is decomposed into simpler, context-preserving phrases using a prompt that includes examples from datasets like MinWikiSplit. This decomposition is dynamic and structure-aware, producing shorter phrases that are easier for LLMs to handle. Next, each phrase is translated independently in a few-shot setting, where relevant examples are retrieved through similarity search methods like BM25. Incorrectly translated phrases in other languages are filtered using a language identifier to maintain quality. Finally, the LLM is prompted again with the original sentence and the self-generated phrase-translation pairs to produce a complete, coherent translation. This approach takes advantage of the fact that LLMs perform better on shorter inputs and can more easily match relevant in-context examples for them. CompTra consistently outperforms traditional few-shot and zero-shot methods across multiple benchmarks, showing strong improvements in low-resource and out-of-domain scenarios.
@@ 4.3.1 Supervised Fine-Tuning (SFT)
Supervised fine-tuning (SFT) involves training an LLM on a large parallel corpus by minimizing the cross-entropy loss between predicted translations and ground truth references. This method significantly enhances the model’s translation accuracy, especially for specialized domains. It has been shown that SFT enables LLMs to achieve strong multilingual translation performance with as few as 32 examples, where training on one language direction generalizes to others [78]. However, using English as the target language can lead to misinterpretations, and the presence of target-side noise tends to affect well-represented languages more adversely. Ensuring successful SFT alignment requires the model to maintain a surface-level focus, preventing it from acquiring unwanted biases during the fine-tuning process. Similarly, fine-tuning improved the translation quality, and generally, large models are known to have better quality output [79]. However, while fine-tuning on parallel data significantly improves overall translation quality as measured by metrics such as COMET, it simultaneously leads to a degradation of several valuable LLM capabilities. Specifically, formality steering, few-shot domain adaptation, and document-level contextualization are all weakened as a result of continued fine-tuning. The authors observe that even with as few as 18,000 parallel training examples, these abilities begin to deteriorate, and the effect intensifies with larger datasets. These findings suggests that fine-tuning must be approached with caution.
@@ 4.3.2 Low-Rank Adaptation (LoRA)
LoRA [80] is a parameter-efficient fine-tuning technique that reduces memory overhead by injecting small trainable matrices into frozen pre-trained models. This approach enables effective adaptation to new translation domains without modifying the entire model. A comprehensive empirical study on parameter-efficient fine-tuning (PEFT) methods for MT, conducted across various parameter budgets and language pairs, observed that using adapters achieves performance on par with full fine-tuning while training only 10% of the parameters [81]. However, the effectiveness of PEFT methods varies significantly based on language pair distance and dataset size, with noticeable struggles for distant language pairs. We refer the readers to [82] and [83] for a detailed analysis for PEFT methods.
@@
An evaluation of 15 publicly available language models on MT tasks showed that QLoRA fine-tuning outperforms both few-shot learning and models trained from scratch, achieving a significant BLEU score improvement while training on only 0.77% of the model parameters [87]. Adapter-based fine-tuning with LoRA was further shown to match the performance of traditional fine-tuning, while reducing the number of training parameters by a factor of 50 [54]. One notable application of PEFT is Language-Specific Fine-Tuning with LoRA (LSFTL) [88], which aims to improve translation quality for low-resource language pairs while maintaining computational efficiency. LSFTL selectively inserts LoRA adapters into the multi-head attention and feed-forward layers of a multilingual Transformer-based MT model. By tuning only a small number of additional parameters it avoids the high cost of full model fine-tuning. Remarkably, the approach allows models like NLLB-200-Distilled-600M to match the performance of much larger models such as NLLB-200-1.3B without LSFTL. Moreover, LSFTL shows faster convergence and better early-stage performance than full-volume fine-tuning.
4.3.3 Monolingual and Parallel data fine-tuning
A two-stage fine-tuning approach has been introduced to enhance MT capabilities in LLMs [89]. Since many LLMs are predominantly trained on English data, the first stage, known as ALMA, fine-tunes the model on non-English monolingual datasets. This step strengthens the model’s proficiency in various languages, enabling a more robust understanding of linguistic nuances across different tongues.
@@ 4.3.3 Monolingual and Parallel data fine-tuning
To address the degradation of LLM capabilities during fine-tuning (Section 4.3.1), the authors in [79] propose a mixed-data fine-tuning strategy that incorporates both monolingual and parallel data. Their experiments demonstrate that this approach not only maintains emergent LLM capabilities but also enhances overall translation quality. For example, formality steering accuracy is preserved more effectively when monolingual data is included, and document-level contextualization shows less decline or even slight improvements compared to parallel-only fine-tuning. This finding suggests that preserving LLM strengths is not inherently at odds with achieving high translation quality. Instead, it points to the need for carefully balanced training objectives that reflect both general translation accuracy and task-specific behaviors. The use of monolingual data—especially data that structurally resembles LLM pretraining corpora—serves as an effective regularizer during adaptation, helping to retain pre-existing generalization capabilities.
@@ 4.3.4 Reinforcement Learning based Finetuning
The standard SFT methodology, which focuses on reducing the gap between predicted outputs and gold-standard references, inherently limits model performance to the quality of the training data. This presents a notable challenge, as even human-authored data—typically regarded as high-quality—can contain imperfections. To address this limitation, the authors in [90] introduced Contrastive Preference Optimization (CPO) for MT. CPO differentiates itself by training models to distinguish between “adequate” and “near-perfect” translations. This is achieved through a contrastive learning strategy that presents the model with hard negative examples—translations that are high-quality but not flawless. By learning to reject these near-perfect yet imperfect translations, the model is encouraged to produce superior outputs. When applied to ALMA models with only 22K parallel sentences and fine-tuning only 0.1% of the parameters, CPO yielded significant improvements. The authors in [91] uses Reinforcement Learning with Human Feedback (RLHF) to improve MT by aligning model outputs with human preferences. Instead of relying on expensive human-annotated preference data, the authors use high-quality human translations from books as a proxy for preferred outputs. The process involves fine-tuning a base translation model on parallel corpora, training a reward model to distinguish between machine and human translations, and then refining the translation model through RL using Proximal Policy Optimization (PPO) [92]. This approach enhances translation quality and shows cross-lingual transferability, meaning the improvements generalize to languages not directly optimized with RLHF. The results demonstrate that RLHF can effectively capture human translation preferences in a cost-efficient manner. Alternative RL-based LLM fine-tuning for MT is to used a fine-grained token-level reward [93]. xCOMET [94] identifies error spans and classifies their severity as minor, major, or critical. Then, each token in an error span is assigned a reward according to a custom severity mapping. To integrate these signals, the authors adapt standard RL algorithms into token-level variants, enabling the model to update policies at each token step rather than after full sequence generation. These token-level RL objectives provide continuous, granular feedback, improving stability and mitigating overfitting and exposure bias. The method shows significant gains across metrics when compared to sentence-level RL and supervised fine-tuning baselines. The authors in [95] propose a rule-metric mixed reward system named MT-R1-Zero that integrates both formatting constraints and automatic quality metrics to guide training. This reward combines a format reward, which enforces structured outputs using think and translate tags, with a metric reward based on established translation quality metrics — BLEU for lexical accuracy, COMETKiwi for semantic fidelity, and their combination for balanced optimization. The training leverages the GRPO algorithm, which computes advantages from multiple candidate translations to stabilize learning. MT-R1-Zero produces high-quality translation outputs, achieving strong performance across in-domain, out-of-distribution, and multilingual settings. Notably, ablation studies confirm that most of the performance gains stem from the RL framework itself, even in the absence of explicit reasoning prompts, highlighting the efficacy of the rule-guided, reward-driven optimization process.
5 LLMs and Encoder-Decoder across Languages